Given a cube with a volume of 27 cm3, what is the volume of a square pyramid that can fit perfectly inside the cube?


3cm3

6cm3

9cm3

12cm3


1
Expert's answer
2020-07-02T19:02:47-0400

Let us determine the edge "a" of the cube. The volume is "V=a^3=27," therefore "a = \sqrt[3]{27} = 3\,\mathrm{cm}."

The square pyramid has the square base of area "S=a^2 = 9\,\mathrm{cm}^2" and the height of "a" . The volume of a pyramid is "V_p = \dfrac13\cdot S\cdot h = \dfrac13\cdot 9\,\mathrm{cm}^2\cdot 3\,\mathrm{cm} = 9\,\mathrm{cm}^3."